OpenStudy (anonymous):

METAL&FAN. ;) What is the value of the coefficient "b" when the quadratic equation y = (2x + 1)(4x − 3) is written in standard form? −2 −3 5 10

OpenStudy (anonymous):

first thing you want to do is multiply it out

OpenStudy (anonymous):

y=8x^2-2x-3

OpenStudy (anonymous):

so b=-2

OpenStudy (anonymous):

The standard form is ax² + bx + c Expend the expression (2x+1)(4x-3) ( Foil ) = 8x² -6x +4x -3 = 8x² -2x -3 b = -3 == i got this. What did i do wrong?

OpenStudy (anonymous):

@mirrorimagemirage

OpenStudy (anonymous):

if the standard form is ax^2+bx+c, then look for the b in your new equation. hint: b is in front of x

OpenStudy (anonymous):

-3 would be the c value

OpenStudy (anonymous):

Oooh thankyou(:

OpenStudy (anonymous):

yeah no problem!